From GlueXWiki
Jump to: navigation, search

March 14, 2013 FDC meeting


  1. Production Construction Tracking (Dave)
  2. Installation schedule (last week), schedule (this week)(Lubomir)
  3. Engineering (Bill)
  4. Electronics (Chris, Nick)
  5. Tests at EEL126 FDC E-log (Beni)
  6. Other


Participants: Bill, Chris, Nick, Simon, Beni, and Lubomir.


- Dave (not at the meeting) prepared the Production Tracking: all cathodes for the spare package are ready. We will built also one extra cathode that is tensioned now. Wire planes: Chris is coming tomorrow to Blue Crab to finish with the population of the last two wire planes. Then it will remain four wire planes (one will be extra spare) to be deadened and we are done with the spare parts.

- On Friday package #1 was moved from 126 to Blue Crab and then package #2 - from Blue Crab to 126. After connecting #1 to Nitrogen it was found cell #5 is leaking; turned out the barb connector for the guage tube was loosely connected. These connectors are glued initially with super glue, but later on some of them (and not on this one) green epoxy was added. Casey was able to glue a new barb connector (not easy to reach the spacer ring since you have only 5mm between the cathode and the wire frame) and fixed the problem. Then green epoxy was added on all the connectors of packages #1 and #4 that didn't have it before (5 out of 12). Bill wants to try with Casey using shoe glue that better sticks to polyethylene.

- Package #3 repair: on Tuesday hubs were installed on the top side and the package put upside down on the installation table. To access cell #6 we had to remove only the end window and the bottom cathode (type-3). We found 1-2mm hair/lint on a field wire in HV sector 2, exactly where the HV problem was. After removing it we closed the package and started flushing it. It took only 4 people for two hours for this operation, but placing back the ground clips, pre-amps, cooling tube (all had to be removed) and restoring the HV connections will take much longer. The plan is to start testing the package tomorrow.

- Pictures on the copper side of the cathode type-3 from package #3 are linked above. There's again some discoloration in the area that sits against the G10 of the wire frame. We had this when keeping the cathode moist over night when we started the production. One can see the G10 fiber pattern, so the explanation could be that some moisture got trapped there, most likely when there were AC problems in 126; this package was there for many months before the two new AC units were installed. From previous experience we know the copper is still conductive even it changes color.


- This week there were significant changes in the re-baselined schedule. The last- and this-week-schedule related to the FDC installation are linked above. Almost all of the FDC installation activities are delayed by 1-2 months, so the insertion into the magnet happens in the beginning of January 2014. From Eugene we know the big change is a result of a manpower spike that had to be smooth out. Beni: if this is the case probably additional manpower from the universities could help. However, Beni learned from Fernando the real reason is the the timing of the receiving of the fADC125 and F1TDC, and their testing/installation in the crates and racks and here students could not help. Expect more discussions in the coming meetings.

- In this case the chance that we will have to leave Blue Crab but the infrastructure in the Hall will not be ready for the FDC installation, is very big. According to Rusty Sprouse the plan for Blue Crab is to extend the lease by one month (till the end of September), but has not been done yet. In the best base we will have to leave end of August, but the installation will start in October according to the new schedule, so we need to find a storage space. ESB has climate control but is very dirty. Bill: what about space in TDEF? Beni: can we move the clean room (or part of it) from Blue Crab to the ESB?


- Bill and Casey finished the manifold installation on package #4. Bill explained details of the procedure using the pictures linked above. Next step is to do tests with pressured gas at Blue Crab. For the rest of the packages, Bill has to do some modifications on the plastic brackets that are used at the 210deg manifold.

Tests at EEL126

- Beni showed HV and threshold scan of package #1 (entry 107 in the FDC E-log). There were some problems in some of the runs at lower HV, but in general the plateau starts at 2050V for 5V threshold and at 2100V at 7V. Beni went up to 2200V and the HV was stable.

- Package #2: Beni already tested the strips of four cells and all the wire channels (except few with noisy electronics) and all channels were OK.